OpenStudy (anonymous):

An ellipse has foci (0, ±3) and vertices (0, ±4). What is the eccentricity of the ellipse? 0 0.25 0.75 1.00

OpenStudy (anonymous):

e=c/a

OpenStudy (anonymous):

c = 3 a = 4 e = 3/4 = 0.75

OpenStudy (anonymous):

its that simple!?

OpenStudy (anonymous):

Yeah...just use the Formula

OpenStudy (anonymous):

okay thank you for the help! I really appreciate it!

OpenStudy (anonymous):

WELCOME